A 25 foot rope is cut into two pieces. One is 5 1/4 longer than the other one. How Long is the shorter piece? HELP
Answer by stanbon(75887)

A 25 foot rope is cut into two pieces. One is 5 1/4 longer than the other one. How Long is the shorter piece? HELP
------
shorter piece: x ft
longer piece: x + (21/4) ft
-------
Equation:
x + x +(21/4) = 25
----
2x = (100/4)-(21/4)
2x = 79/4
x = 79/8
x = 9 7/8 ft
==============================
Cheers,
Stan H.
================